OpenStudy (mendicant_bias):

Another IVP; am trying to come to the conclusion independently but can already see some issues, will post my workings as I go through the problem.

OpenStudy (mendicant_bias):

\[y''-y'=e^tcos(t), \ \ \ y(0)=0, \ y'(0)=4\]

OpenStudy (anonymous):

doing good

OpenStudy (mendicant_bias):

\[s^2Y(s)-sy(0)-y'(0)-[sY(s)-y(0)]= \frac{1}{(s-1)^2+1}\]

OpenStudy (mendicant_bias):

\[Y(s)[s^2-s]-4=\frac{1}{(s-1)^2+1}\]

OpenStudy (mendicant_bias):

Now here's where I can see things getting really bad in the denominator of the fractions that would result when I isolate Y(s); is this all correct so far, and is it really supposed to look this way, lol?

OpenStudy (anonymous):

Seems fine to me so far.

OpenStudy (mendicant_bias):

\[Y(s)=\frac{1}{[(s-1)^2+1](s^2-s)}+\frac{4}{s^2-s}\]

OpenStudy (mendicant_bias):

Is it a good idea to expand the denominator and then try to simplify it in a different form here, or what should be the next step, using only t or x-axis shifting methods (exponential) to change this? No Heaviside, no convolution, etc.

OpenStudy (mendicant_bias):

\[\frac{1}{[(s^2-2s+1)+1](s^2-s)}+\frac{4}{s^2-s}=\frac{1}{(s^2-2s+2)(s^2-s)}+\frac{4}{s^2-s}\]

OpenStudy (sidsiddhartha):

ok u want to use any other properties?

OpenStudy (sidsiddhartha):

other than partial fractions

OpenStudy (mendicant_bias):

Only allowed to use those shifting ones xP

OpenStudy (mendicant_bias):

Oh, if you mean algebraic ones, then yeah

OpenStudy (anonymous):

I don't see why you cant use partial fractions. Thats pretty much what laplace transform problems are, exercises in partial fractions, lol.

OpenStudy (mendicant_bias):

Can't use Heaviside, can't use Convolution, just reminding; purely algebraic properties to get it into a preferable form, if you see one.

OpenStudy (sidsiddhartha):

ok \[F(s)=\frac{ 1 }{ s(s-1)[(s-1)^2+4] }+\frac{ 4 }{ s(s-1) }\\=[\frac{ 1 }{ s-1 }-\frac{ 1 }{ s }][\frac{ 1 }{ (s-1)^2+4 }]+4[\frac{ 1 }{ s-1 }-\frac{ 1 }{ s }]\]

OpenStudy (sidsiddhartha):

\[=\frac{ 1 }{ s-1 }*\frac{ 1 }{ (s-1)^2+4 }-\frac{ 1 }{ s }*\frac{ 1 }{ (s-1)^2+4 }+\frac{ 4 }{ s-1 }-\frac{ 4 }{ s }\] getting this?

OpenStudy (mendicant_bias):

Man, Siddhartha, you are like a math God, I can't even understand how you're able to figure this stuff out so quickly half of the time. I don't understand what you just did, but I'm going to keep looking at it for a minute. The first post confused me.

OpenStudy (mendicant_bias):

Alright, I'm seeing it, however

OpenStudy (sidsiddhartha):

i just broke \[\frac{ 1 }{ s(s-1) }=\frac{ 1 }{ s-1 }-\frac{ 1 }{ s }\] in the first step

OpenStudy (mendicant_bias):

I'm not immediately seeing how the four is turning up in the denominator of the first fraction, yeah one sec:

OpenStudy (sidsiddhartha):

because u gotta use some shifting property

OpenStudy (mendicant_bias):

Wait whoah

OpenStudy (mendicant_bias):

Are you using convolution in there later on?

OpenStudy (sidsiddhartha):

now first only taking this term \[\frac{ 1 }{ s-1 }*[\frac{ 1 }{ (s-1)^2+4 }]\] dont worry i wont use conv

OpenStudy (mendicant_bias):

I just don't get where that four came from in the denominator, in the very beginning.

OpenStudy (anonymous):

\[\frac{ 1 }{ [(s-1)^{2}+1](s^{2}-s) }+ \frac{ 4 }{ s^{2} - s } = \frac{ -s }{ (s-1)^{2} + 1 } + \frac{ 5 }{ s-1 }- \frac{ 9 }{ 2s }\] This is another partial fraction result if this is more comfortable looking to ya.

OpenStudy (sidsiddhartha):

so first use shifting on this one\[e^{t}*L^{-1}[\frac{ 1 }{ s }*\frac{ 1 }{ s^2+4 }]\]

OpenStudy (mendicant_bias):

\[[(s-1)^2+1](s^2-s) = s(s-1)[(s-1)^2+1]\]

OpenStudy (mendicant_bias):

That's what I see happening to the denominator in the beginning, not sure where that four came from

OpenStudy (sidsiddhartha):

i know @concentrationalizing that partial fraction will work ,but i'm just trying some thing else to solve it :)

OpenStudy (anonymous):

Yeah, Im curious how you're messing with it, but alright, lol.

OpenStudy (sidsiddhartha):

u, sorry that 4 wont be there its a typo

OpenStudy (sidsiddhartha):

yes so lets deal with the first term,\[\frac{ 1 }{ s-1 }*\frac{ 1 }{ (s-1)^2+1 }\] thats it right?

OpenStudy (mendicant_bias):

I'm just confused that you are using the sign typically used for convolution for multiplication; that's what's going on, yes? And-oh, okay, the four is a typo, I think I can follow, let me reread everything with that in mind

OpenStudy (sidsiddhartha):

ok sorry it not conv simple multiplication

OpenStudy (mendicant_bias):

When you did\[\frac{1}{s(s-1)}=\frac{1}{s-1}-\frac{1}{s}\], that was partial fractions in the first place, yes? Or was that something else?

OpenStudy (sidsiddhartha):

yes a little one :P

OpenStudy (anonymous):

That part is partial fractions, yes. The rest he seems to be making fancy, lol.

OpenStudy (mendicant_bias):

Alright, then one sec, rereading with that in knowledge. Just filling in the gaps.

OpenStudy (sidsiddhartha):

yes now i have just multiplied this with terms in the right hand side

OpenStudy (mendicant_bias):

Just writing out the math myself, give me a moment.

OpenStudy (sidsiddhartha):

ok take the first term in RHs\[[\frac{ 1 }{ s-1 }-\frac{ 1 }{ s }].[\frac{ 1 }{ (s-1)^2+1 }]\\=\frac{ 1 }{ s-1 }.\frac{ 1 }{ (s-1)^2 +1}-\frac{ 1 }{ s }.\frac{ 1 }{ (s-1)^2+1 }\]

OpenStudy (mendicant_bias):

\[\frac{1}{[(s-1)^2+1](s^2-s)}-\frac{4}{(s^2-s)}\]\[\frac{1}{s(s-1)[(s-1)^2+1]}-\frac{4}{s(s-1)}=\left\{\vphantom{}\right\}\] It's 1 AM, I think I'm going to sleep for a little and then wake back up and work on this-sorry for suddenly leaving.

OpenStudy (sidsiddhartha):

ok sleep well but beware laplace will haunt you in your dreams bwahahaha

OpenStudy (sidsiddhartha):

lol :)

OpenStudy (mendicant_bias):

Nevermind, can't sleep like this anyways, finishing this off.

OpenStudy (mendicant_bias):

\[\frac{1}{s(s-1)[(s-1)^2+1]}-\frac{4}{s(s-1)}=\]\[\Bigg[\frac{1}{s-1}-\frac{1}{s}\Bigg]\Bigg[\frac{1}{(s-1)^2+1}\Bigg]-4\Bigg[\frac{1}{s-1}-\frac{1}{s}\Bigg]\]

OpenStudy (anonymous):

I'll just finish it out the way I was setting it up :) So given that I can split it into partial fractions and get: \[\frac{ -s }{ (s-1)^{2} + 1 }+ \frac{ 5 }{ s-1 } - \frac{ 9 }{ 2s }\] So the last two fractions don't require any fancy manipulation at all. we just have \(\frac{5}{s-1}\) matching the form \(F(s) = \frac{1}{s-a} = e^{at}\) and we have \(\frac{-9}{2s}\) matching the form \(F(s) = \frac{1}{s} = 1\). Of course the 1's in the F(s) formulas will be replace with our actual coefficients. The first fraction is close to a form that has a nice inverse transform, we just have to add something to it. So I'll do this: \[\frac{ -s }{ (s-1)^{2} + 1 } = \frac{ -s+1-1 }{ (s-1)^{2} + 1 } = \frac{ -(s-1) - 1 }{ (s-1)^{2} + 1 } = -\frac{ (s-1) }{ (s-1)^{2} + 1 } - \frac{ 1 }{ (s-1)^{2} + 1 }\] So splitting this up gives me the forms: \(F(s) = \frac{s-a}{(s-a)^{2} + k^{2}} = e^{at}cos kt\) and \(F(s) = \frac{k}{(s-a)^{2} + k^{2}} = e^{at}sin kt\) So with that, I can do the inverse laplace transform now to all of our fractions and get: \[-L^{-1}[\frac{ (s-1) }{ (s-1)^{2} + 1 }] - L^{-1}[\frac{ 1 }{ (s-1)^{2} + 1 }] + L^{-1}[\frac{ 5 }{ s-1 }]-L^{-1}[\frac{ 9 }{ 2s }]\] \[= -e^{t}cost - e^{t}sint + 5e^{t} - \frac{ 9 }{ 2 }\] Just my way of doing it, though.

OpenStudy (sidsiddhartha):

yes that seems ok from here u can multiply then use some shifting properties ,

OpenStudy (sidsiddhartha):

but i think this will be more difficult because at some point u have to use some integral because of that (1/s) term \[1/s[f(s)]=\int\limits_{0}^{t}f(t)dt\] thats why partial fraction is more suitable i think :)

OpenStudy (sidsiddhartha):

sorry, i'll have to head out for college now

OpenStudy (sidsiddhartha):

i'll post the solution when i'll get back :)

OpenStudy (mendicant_bias):

\[\Bigg[\frac{1}{(s-1)[(s-1)^2+1]}\Bigg]-\Bigg[\frac{1}{s[(s-1)^2+1]}\Bigg]-4\Bigg[\frac{1}{s-1}-\frac{1}{s}\Bigg]\]

OpenStudy (sidsiddhartha):

yes that fine :)

OpenStudy (mendicant_bias):

Typo, plus sign supposed to be there, my bad

OpenStudy (mendicant_bias):

Alright, now next

OpenStudy (sidsiddhartha):

next use shifting property, and sorry i gotta go now :(

OpenStudy (mendicant_bias):

Sure thing! Ttyl, thank you, both you guys.

OpenStudy (anonymous):

Im still randomly around if you wanna mess with it more, lol. Seems like you're more along the lines of sid's approach, but I can still watch on and keep track of your steps and such.

OpenStudy (mendicant_bias):

I'm along the line of any approach, I'd like to try both, just sticking with that one and finishing it before the other, heh.

OpenStudy (mendicant_bias):

(Just thinking about how to move forward, gimme a minute)

OpenStudy (mendicant_bias):

I dunno, shifting aside, I've never really dealt with inverse Laplace Transforms of this general form; how should I move forward?

OpenStudy (anonymous):

Well, when sid starting splitting it up as he did, I thought he was going along the lines of a convolution as well. That's why it's hard for me to see this being done without further partial fractions.